# Theory of Photocurrent and High-Harmonic Generation with Chiral Fermions

### v1

@@ Abstract
Introduction.—The nonlinear optical response induced by high-intensity lasers is a central topic in modern condensed matter physics Bloembergen (1996); Boyd (2008); Nagaosa (2022); Morimoto et al. (2023). In particular, understanding the mechanisms of the bulk photovoltaic effect and high-harmonic generation (HHG) is of great interest. The bulk photovoltaic effect generates dc photocurrent under irradiation and is allowed in noncentrosymmetric materials von Baltz and Kraut (1981); Sturman and Fridkin (1992); Sipe and Shkrebtii (2000); Young and Rappe (2012); Young et al. (2012); Morimoto and Nagaosa (2016); Dai and Rappe (2023). Recently, dc photocurrent generation in noncentrosymmetric Weyl semimetals has attracted significant attention Ma et al. (2017); Sun et al. (2017); Osterhoudt et al. (2019); Sirica et al. (2019); Wang et al. (2019); Ma et al. (2019); Gao et al. (2020). Additionally, since the first observation of HHG in semiconductors Ghimire et al. (2011), HHG has been extensively studied in various solids Ghimire and Reis (2019); Yue and Gaarde (2022); Li et al. (2023); Bhattacharya et al. (2023); Hirori et al. (2024). HHG has been observed in topological materials, such as Weyl semimetals Patankar et al. (2018); Lv et al. (2021), Dirac semimetals Cheng et al. (2020); Kovalev et al. (2020), and topological insulators Bai et al. (2021); Schmid et al. (2021); Baykusheva et al. (2021). These studies contribute to the dynamic control of chiral fermions in solids and the exploration of quantum phenomena in strongly driven systems Oka and Aoki (2009); Oka and Kitamura (2019); Rudner and Lindner (2020).
B20-type chiral semimetals have long been studied in various contexts, for example, as thermoelectric materials Asanabe et al. (1964); McNeill and Ware (1964); Rowe (1995), and their significance as a platform for hosting unconventional chiral fermions Bradlyn et al. (2016); Rarita and Schwinger (1941); Mañes (2012); Wieder et al. (2016); Weng et al. (2016a, b); Zhu et al. (2016); Liang and Yu (2016); Ezawa (2016) has recently attracted growing interest. Several theoretical studies have predicted the emergence of chiral multifold fermions in B20-type chiral semimetals, such as CoSi Tang et al. (2017); Pshenay-Severin et al. (2018) and RhSi Chang et al. (2017), and recent experiments have demonstrated characteristic features of the chiral multifold fermions through optical responses Xu et al. (2020); Maulana et al. (2020) and angle-resolved photoemission spectroscopy Lv et al. (2017); Takane et al. (2019); Sanchez et al. (2019); Schröter et al. (2019); Rao et al. (2019). Moreover, the circular photogalvanic effect, a hallmark of nonlinear optical responses with chiral fermions, has been observed Rees et al. (2020); Ni et al. (2020, 2021). It is a natural extension of interest to explore characteristic nonlinear optical responses of B20-type chiral semimetals induced by intense laser irradiation, where nonequilibrium effects become more pronounced. Such investigations remain a promising avenue for further exploration Ezawa (2017); Ahn and Ghosh (2023); Fan et al. (2024).
@@
In this Letter, we propose that bulk harmonic photocurrent of arbitrary order, consisting of both dc and ac components, is generated in B20-type chiral semimetals by circularly polarized light irradiation. We consider CoSi irradiated by circularly polarized light with electric field $E$ and frequency $\Omega$, as shown in Fig. 1(a). The dc-photocurrent generation corresponding to the circular photogalvanic effect manifests as a nonlinear optical response. The dc-photocurrent exhibits multiple sign changes as a function of light parameters, $E$ and $\Omega$. Additionally, the ac harmonic photocurrent gives rise to HHG. The HHG intensity exhibits parity-dependent directivity, which means odd- and even-order harmonics are directed perpendicular and parallel to the polarization plane, respectively. The directivity originates from the combination of the time-periodic driving by light and the lattice symmetries of chiral semimetals. These findings advance the understanding of nonlinear optical responses in chiral systems and open new avenues for next-generation optoelectronic applications.
Formalism.— We start with a four-band tight-binding model to describe the electronic structure of CoSi Chang et al. (2017); Flicker et al. (2018). The model qualitatively reproduces the band structure of CoSi near the Fermi level and provides a reasonable starting point for capturing the essential features of B20-type chiral semimetals. The model Hamiltonian is given by
@@
where $c_{jX}^{(\dagger)}$ denotes the annihilation (creation) operator of an electron at the $X$ sublattice of the $j$th site, $h_{jX,kY}$ represents a transfer integral, and ${\bf\it\delta}_{jX,kY}$ represents the vector from the $Y$ sublattice of the $k$th site to the $X$ sublattice of the $j$th site. The coordinates of each sublattice are $(0,0,0)$, $(a/2,a/2,0)$, $(a/2,0,a/2)$, and $(0,a/2,a/2)$, where the lattice constant is $a=0.44~{}\mathrm{nm}$. The effect of the light irradiation is included through the Peierls phase. The vector potential is given by ${\bf\it A}(t)=(E/\Omega)(\sin\Omega t,\cos\Omega t,0)$, where $E$ and $\Omega$ are the amplitude and frequency of the light electric field, respectively. In the above model, spin-orbit coupling is neglected, and the spin degree of freedom simply doubles the photocurrent.
We consider the nearest-neighbor and next-nearest-neighbor hopping processes, where their transfer integrals are schematically illustrated in Fig. 1(b). The nearest-neighbor transfer integral is given by $h_{\pm}=h_{1}\pm h_{c}$, where the sign depends on the bond. The blue and orange bonds correspond to the positive and negative sign, respectively. A finite $h_{\mathrm{c}}$ introduces chirality into the model. The next-nearest-neighbor transfer integral is given by $h_{2}$. The transfer integrals are set as $h_{1}=0.34~{}\mathrm{eV}$, $h_{\mathrm{c}}=0.14~{}\mathrm{eV}$, and $h_{2}=0.13~{}\mathrm{eV}$ to reproduce DFT bands of CoSi. Additionally, the on-site energy is set to $h_{0}=0.60~{}\mathrm{eV}$ to appropriately tune the Fermi level. The energy band structure without light irradiation along the path in the Brillouin zone shown in Fig. 1(c) is presented in Fig. 1(d). The spin-1 chiral fermion is observed at the $\Gamma$ point, and the double Weyl fermion at the R point Tang et al. (2017); Chang et al. (2017). The band structure is consistent with the DFT calculations Tang et al. (2017); Xu et al. (2020).
We use the Floquet-Keldysh formalism to describe the nonequilibrium states driven by light irradiation Tsuji et al. (2008, 2009); Aoki et al. (2014). We assume that a time-periodic nonequilibrium state is realized through interactions with a reservoir. The single particle Green’s function is given by $G_{{\bf\it k},XY}(t,t^{\prime})=-(i/\hbar)\langle T_{C}c_{{\bf\it k}X,\mathrm{H}}(t)c_{{\bf\it k}Y,\mathrm{H}}^{\dagger}(t^{\prime})\rangle$. Using the Floquet Green’s function, the Dyson equation is written as
@@
$I_{\parallel}(n)=\absolutevalue{n\Omega J_{\parallel,n}}^{2},~{}I_{\perp}(n)=\absolutevalue{n\Omega J_{\perp,n}}^{2},$ (5)
@@
Figure 2: (a) DC photocurrrent $J_{z,0}$ in plane of $E$ and $\Omega$ (upper panel) and those for various values of $\Omega$ as a function of $E$ (lower panel) for a narrower window of light intensity |$\absolutevalue{E}\leq 1~{}\mathrm{MV/cm}$. Here, the unit is $J_{0}=\frac{2}{(2\pi)^{4}}\frac{e^{2}}{\hbar a^{2}}\mathrm{V}\simeq 1.6\times 10^{12}\mathrm{A/m^{2}}$. Inset shows the coefficient $\beta_{zz}$. (b) Nonequilibrium electron distribution spectrum $N_{{\bf\it k}}(\omega)$ across $\Gamma$ point (upper panel) and that across R point (lower panel) in the $k_{z}$ direction for a relatively weaker light intensity of $E=1~{}\mathrm{MV/cm}$. The dashed curves indicate the Floquet bands. We set $\gamma=0.05~{}\mathrm{eV}$ and $\beta^{-1}=0.03~{}\mathrm{eV}$.
@@
Figure 3: (a) DC photocurrrent $J_{z,0}$ in plane of $E$ and $\Omega$ (upper panel) and those for various values of $\Omega$ as a function of $E$ (lower panel) for a larger window of light intensity |$\absolutevalue{E}\leq 4~{}\mathrm{MV/cm}$. The unit is $J_{0}=\frac{2}{(2\pi)^{4}}\frac{e^{2}}{\hbar a^{2}}\mathrm{V}\simeq 1.6\times 10^{12}\mathrm{A/m^{2}}$. (b) Nonequilibrium electron distribution spectrum $N_{{\bf\it k}}(\omega)$ across $\Gamma$ point (upper panel) and that across R point (lower panel) in the $k_{z}$ direction with for a relatively stronger light intensity of $E=4~{}\mathrm{MV/cm}$. The dashed curves indicate the Floquet bands. We set $\gamma=0.05~{}\mathrm{eV}$ and $\beta^{-1}=0.03~{}\mathrm{eV}$.
@@
Figure 4: (a) HHG intensity $I_{\parallel}$ and $I_{\perp}$ as a function of $n$ for several values of $E$. The unit is given by $I_{0}=(J_{0}\hbar^{-1}~{}\mathrm{eV})^{2}$. (b) $I_{\parallel}$ and $I_{\perp}$ as functions of $E$ for odd and even $n$, respectively. (c) $I_{\parallel}$ and $I_{\perp}$ as functions of $\hbar\Omega$ for odd and even $n$, respectively. We set $\gamma=0.05~{}\mathrm{eV}$ and $\beta^{-1}=0.03~{}\mathrm{eV}$.
@@
Results.—In our setup, both the dc photocurrent and HHG arise. First, we show the results of the dc photocurrent in Figs. 2 and 3, and then show the results of HHG intensity in Fig. 4. The photon number is restricted to $|m|\leq 10$ in all numerical calculations, and Eq. (3) is discretized and numerically integrated over wavenumber ${\bf\it k}$ and frequency $\omega$ using a $40^{3}\times 60$ mesh. Under these conditions, the numerical results show sufficient convergence, ensuring that the qualitative features remain unchanged and further increasing the mesh resolution leads to only minor quantitative corrections.
@@
Figure 2(a) shows the dc photocurrent $J_{z,0}$ as a function of the electric field $E$ and the frequency $\Omega$. The bottom panel shows the corresponding $J_{z,0}$ curves for specific values of $\Omega$. For a weak $E$-field ($|E|\ll 1~{}\mathrm{MV/cm}$), the photocurrent exhibits a quadratic dependence on $E$, described as $J_{z,0}=\beta_{zz}\tau E^{2}$, where $\tau=\hbar/2\gamma$ is the relaxation time. The coefficient $\beta_{zz}$ is expected to be quantized under appropriate conditions de Juan et al. (2017); Flicker et al. (2018); Orenstein et al. (2021). Here, we evaluate the coefficient $\beta_{zz}$ by fitting the calculated photocurrent $J_{z,0}$ in the weak $E$-field regime to a quadratic curve. The coefficient $\beta_{zz}$ is shown in the inset of the bottom panel. $\beta_{zz}$ changes its sign from negative to positive around $\hbar\Omega=0.3~{}\mathrm{eV}$, and reaches a maximum near $\hbar\Omega=0.4~{}\mathrm{eV}$. Under the current conditions, the quantization of $\beta_{zz}$ is not observed, consistent with previous numerical simulations Ni et al. (2020, 2021). As the $E$-field increases, higher-order corrections become significant, leading to deviations from the quadratic approximation. Above $E\simeq 0.5~{}\mathrm{MV/cm}$, $J_{z,0}$ as a function of $\Omega$ exhibits multiple sign changes, where higher-order terms in powers of $E$ play a significant role in determining the sign of $J_{z,0}$. Note that the sign of $J_{z,0}$ can be reversed by changing the handedness of circularly polarized light, reflecting the chirality of CoSi.
@@
Figure 3(a) shows the photocurrent $J_{z,0}$, including the strong $E$-field regime $(|E|\gtrsim 1~{}\mathrm{MV/cm})$. At relatively low frequencies $\Omega\lesssim 0.5~{}\mathrm{eV}$, $J_{z,0}$ is positive and increases with the $E$-field, whereas at higher frequencies $\Omega\gtrsim 0.5~{}\mathrm{eV}$, $J_{z,0}$ becomes negative. In this regime, the nonlinearity becomes more pronounced, and the electronic structure deviates significantly from equilibrium. Figure 3(b) shows the spectrum $N_{{\bf\it k}\omega}$ under the intense $E$-field condition. The spectrum reveal highly asymmetric features, reflecting the hybridization of photon-dressed bands. The spectrum is broadened and widely distributed, further reflecting the strongly driven nonequilibrium nature of the chiral system.
@@
The parity-dependent directivity of HHG can be understood based on dynamical symmetry Tang and Rabin (1971); Alon et al. (1998); Saito et al. (2017); Neufeld et al. (2019); Ikeda and Sato (2019); Ikeda (2020); Kanega et al. (2021); Kanega and Sato (2024). From the combination of the screw rotation symmetry of B20-type materials, $s_{2z}=\{C_{2z}|\frac{1}{2}0\frac{1}{2}\}$, and the time periodicity of the photocurrent $J_{\alpha}(t+T)=J_{\alpha}(t)$, we obtain $J_{\parallel}(t+T/2)=e^{-i\pi}J_{\parallel}(t)$ and $J_{\perp}(t+T/2)=J_{\perp}(t)$. These relations result in the following selection rules for HHG: $I_{\parallel}(n)=0$ except for $n\equiv 1\pmod{2}$ and $I_{\perp}(n)=0$ except for $n\equiv 0\pmod{2}$.
To clarify the $E$-field dependence of HHG in more detail, Figure 4(b) shows log-log plots of the HHG intensity as a function of $E$. In the region where the $E$-field is relatively weak ($E\lesssim 1~{}\mathrm{MV/cm}$), the $n$th order HHG intensity is proportional to $E^{2n}$. This behavior is consistent with the $n$th order response of $J_{\alpha,n}(\propto E^{n})$ in the perturbative regime. As the $E$-field increases, the $E^{2n}$ lines intersect in $E\gtrsim 1~{}\mathrm{MV/cm}$, indicating the breakdown of perturbative picture. In the strong $E$-field regime, non-perturbative effects become significant. These effects include highly hybridized photon-dressed bands and contributions from high-energy photon-dressed states. As a result, deviations from the $E^{2n}$ lines are observed.
Finally, the frequency $\Omega$ dependence of the HHG intensity is explained. Figure 4(c) shows semilog plots of the HHG intensity as a function of $\Omega$ for the fixed value of $E=2.0~{}\mathrm{MV/cm}$. For $n=1$, $I_{\parallel}$ increases monotonically with $\Omega$. For $n>2$, the HHG intensity tends to decrease as $\Omega$ increases. Furthermore, as $n$ increases, the reduction in HHG intensity with increasing $\Omega$ becomes more pronounced. This suggests that HHG detection is more feasible at lower frequencies.
Discussion.—The selection rules discussed above can be extended to systems with an $N$-fold rotational or screw symmetry axis. When light propagates along either of these axes, the in-plane and out-of-plane photocurrents satisfy the following conditions:

### v2

@@ Abstract
I Introduction
The nonlinear optical response induced by high-intensity lasers is a central topic in modern condensed matter physics [1, 2, 3, 4]. In particular, understanding the mechanisms of the bulk photovoltaic effect and high-harmonic generation (HHG) is of great interest. The bulk photovoltaic effect generates dc photocurrent under irradiation and is allowed in noncentrosymmetric materials [5, 6, 7, 8, 9, 10, 11]. Recently, dc photocurrent generation in noncentrosymmetric Weyl semimetals has attracted significant attention [12, 13, 14, 15, 16, 17, 18]. Additionally, since the first observation of HHG in semiconductors [19], HHG has been extensively studied in various solids [20, 21, 22, 23, 24]. HHG has been observed in topological materials, such as Weyl semimetals [25, 26], Dirac semimetals [27, 28], and topological insulators [29, 30, 31]. These studies contribute to the dynamic control of chiral fermions in solids and the exploration of quantum phenomena in strongly driven systems [32, 33, 34].
B20-type chiral semimetals have long been studied in various contexts, for example, as thermoelectric materials [35, 36, 37], and their significance as a platform for hosting unconventional chiral fermions [38, 39, 40, 41, 42, 43, 44, 45, 46] has recently attracted growing interest. Several theoretical studies have predicted the emergence of chiral multifold fermions in B20-type chiral semimetals, such as CoSi [47, 48] and RhSi [49], and recent experiments have demonstrated characteristic features of the chiral multifold fermions through optical responses [50, 51] and angle-resolved photoemission spectroscopy [52, 53, 54, 55, 56, 57]. Moreover, the circular photogalvanic effect, a hallmark of nonlinear optical responses with chiral fermions, has been observed [58, 59, 60]. It is a natural extension of interest to explore characteristic nonlinear optical responses of B20-type chiral semimetals induced by intense laser irradiation, where nonequilibrium effects become more pronounced. Such investigations remain a promising avenue for further exploration [61, 62, 63].
@@ I Introduction
In this paper, we propose that bulk harmonic photocurrent of arbitrary order, consisting of both dc and ac components, is generated in B20-type chiral semimetals by circularly polarized light irradiation. We consider CoSi irradiated by circularly polarized light with electric field $E$ and frequency $\Omega$, as shown in Fig. 1(a). Light propagates in the (001) direction, which is suitable for the symmetry analysis. The dc-photocurrent generation corresponding to the circular photogalvanic effect manifests as a nonlinear optical response. The dc-photocurrent exhibits multiple sign changes as a function of light parameters, $E$ and $\Omega$. Additionally, the ac harmonic photocurrent gives rise to HHG. The HHG intensity exhibits parity-dependent directivity, which means odd- and even-order harmonics are directed perpendicular and parallel to the polarization plane, respectively. The directivity originates from the combination of the time-periodic driving by light and the lattice symmetries of chiral semimetals. These findings advance the understanding of nonlinear optical responses in chiral systems and open new avenues for next-generation optoelectronic applications.
The paper is organized as follows. In Sec. II, we introduce the model Hamiltonian describing the electronic structure of CoSi under laser driving and the Floquet-Keldysh formalism for calculating the photocurrent. In Sec. III, we present the numerical results and explain the detailed behavior of the dc photocurrent and HHG. In Sec. IV, we briefly discuss the selection rules for HHG, the experimental feasibility of our setup, and the effects of spin-orbit coupling. Our conclusions are given in Sec. V.
II Formalism
We start with a four-band tight-binding model to describe the electronic structure of CoSi [49, 64]. The model qualitatively reproduces the band structure of CoSi near the Fermi level and provides a reasonable starting point for capturing the essential features of B20-type chiral semimetals. The model Hamiltonian is given by
@@ II Formalism
where $c_{jX}^{(\dagger)}$ denotes the annihilation (creation) operator of an electron at the $X$ sublattice of the $j$th site, $h_{jX,kY}$ represents a transfer integral, and ${\bf\it\delta}_{jX,kY}$ represents the vector from the $Y$ sublattice of the $k$th site to the $X$ sublattice of the $j$th site. The coordinates of each sublattice are $(0,0,0)$, $(a/2,a/2,0)$, $(a/2,0,a/2)$, and $(0,a/2,a/2)$, where the lattice constant is $a=0.44\leavevmode\nobreak\ \mathrm{nm}$. The effect of the light irradiation is included through the Peierls phase. The vector potential is given by ${\bf\it A}(t)=(E/\Omega)(\sin\Omega t,\cos\Omega t,0)$, where $E$ and $\Omega$ are the amplitude and frequency of the light electric field, respectively. In the above model, spin-orbit coupling is neglected, and the spin degree of freedom simply doubles the photocurrent.
We consider the nearest-neighbor and next-nearest-neighbor hopping processes, where their transfer integrals are schematically illustrated in Fig. 1(b). The nearest-neighbor transfer integral is given by $h_{\pm}=h_{1}\pm h_{c}$, where the sign depends on the bond. The blue and orange bonds correspond to the positive and negative sign, respectively. A finite $h_{\mathrm{c}}$ introduces chirality into the model. The next-nearest-neighbor transfer integral is given by $h_{2}$. The transfer integrals are set as $h_{1}=0.34\leavevmode\nobreak\ \mathrm{eV}$, $h_{\mathrm{c}}=0.14\leavevmode\nobreak\ \mathrm{eV}$, and $h_{2}=0.13\leavevmode\nobreak\ \mathrm{eV}$ to reproduce DFT bands of CoSi. Additionally, the on-site energy is set to $h_{0}=0.60\leavevmode\nobreak\ \mathrm{eV}$ to appropriately tune the Fermi level. The energy band structure without light irradiation along the path in the Brillouin zone shown in Fig. 1(c) is presented in Fig. 1(d). The spin-1 chiral fermion is observed at the $\Gamma$ point, and the double Weyl fermion at the R point [47, 49]. The band structure is consistent with the DFT calculations [47, 50].
We use the Floquet-Keldysh formalism to describe the nonequilibrium states driven by light irradiation [65, 66, 67]. We assume that a time-periodic nonequilibrium state is realized through interactions with a reservoir. The single particle Green’s function is given by $G_{{\bf\it k},XY}(t,t^{\prime})=-(i/\hbar)\langle T_{C}c_{{\bf\it k}X,\mathrm{H}}(t)c_{{\bf\it k}Y,\mathrm{H}}^{\dagger}(t^{\prime})\rangle$. Using the Floquet Green’s function, the Dyson equation is written as
@@
$I_{\parallel}(n)=\absolutevalue{n\Omega J_{\parallel,n}}^{2},\leavevmode\nobreak\ I_{\perp}(n)=\absolutevalue{n\Omega J_{\perp,n}}^{2},$ (5)
@@
Figure 2: (a) DC photocurrrent $J_{z,0}$ in plane of $E$ and $\Omega$ (upper panel) and those for various values of $\Omega$ as a function of $E$ (lower panel) for a narrower window of light intensity |$\absolutevalue{E}\leq 1\leavevmode\nobreak\ \mathrm{MV/cm}$. Here, the unit is $J_{0}=\frac{2}{(2\pi)^{4}}\frac{e^{2}}{\hbar a^{2}}\mathrm{V}\simeq 1.6\times 10^{12}\mathrm{A/m^{2}}$. Inset shows the coefficient $\beta_{zz}$. (b) Nonequilibrium electron distribution spectrum $N_{{\bf\it k}}(\omega)$ across $\Gamma$ point (upper panel) and that across R point (lower panel) in the $k_{z}$ direction for a relatively weaker light intensity of $E=1\leavevmode\nobreak\ \mathrm{MV/cm}$. The dashed curves indicate the Floquet bands. We set $\gamma=0.05\leavevmode\nobreak\ \mathrm{eV}$ and $\beta^{-1}=0.03\leavevmode\nobreak\ \mathrm{eV}$.
@@
Figure 3: (a) DC photocurrrent $J_{z,0}$ in plane of $E$ and $\Omega$ (upper panel) and those for various values of $\Omega$ as a function of $E$ (lower panel) for a larger window of light intensity |$\absolutevalue{E}\leq 4\leavevmode\nobreak\ \mathrm{MV/cm}$. The unit is $J_{0}=\frac{2}{(2\pi)^{4}}\frac{e^{2}}{\hbar a^{2}}\mathrm{V}\simeq 1.6\times 10^{12}\mathrm{A/m^{2}}$. (b) Nonequilibrium electron distribution spectrum $N_{{\bf\it k}}(\omega)$ across $\Gamma$ point (upper panel) and that across R point (lower panel) in the $k_{z}$ direction with for a relatively stronger light intensity of $E=4\leavevmode\nobreak\ \mathrm{MV/cm}$. The dashed curves indicate the Floquet bands. We set $\gamma=0.05\leavevmode\nobreak\ \mathrm{eV}$ and $\beta^{-1}=0.03\leavevmode\nobreak\ \mathrm{eV}$.
@@
Figure 4: (a) HHG intensity $I_{\parallel}$ and $I_{\perp}$ as a function of $n$ for several values of $E$. The unit is given by $I_{0}=(J_{0}\hbar^{-1}\leavevmode\nobreak\ \mathrm{eV})^{2}$. (b) $I_{\parallel}$ and $I_{\perp}$ as functions of $E$ for odd and even $n$, respectively. (c) $I_{\parallel}$ and $I_{\perp}$ as functions of $\hbar\Omega$ for odd and even $n$, respectively. We set $\gamma=0.05\leavevmode\nobreak\ \mathrm{eV}$ and $\beta^{-1}=0.03\leavevmode\nobreak\ \mathrm{eV}$.
III Results
@@
In our setup, both the dc photocurrent and HHG arise. First, we show the results of the dc photocurrent in Figs. 2 and 3, and then show the results of HHG intensity in Fig. 4. The photon number is restricted to $|m|\leq 10$ in all numerical calculations, and Eq. (3) is discretized and numerically integrated over wavenumber ${\bf\it k}$ and frequency $\omega$ using a $40^{3}\times 60$ mesh. Under these conditions, the numerical results show sufficient convergence, ensuring that the qualitative features remain unchanged and further increasing the mesh resolution leads to only minor quantitative corrections.
@@
Figure 2(a) shows the dc photocurrent $J_{z,0}$ as a function of the electric field $E$ and the frequency $\Omega$. The bottom panel shows the corresponding $J_{z,0}$ curves for specific values of $\Omega$. For a weak $E$-field ($|E|\ll 1\leavevmode\nobreak\ \mathrm{MV/cm}$), the photocurrent exhibits a quadratic dependence on $E$, described as $J_{z,0}=\beta_{zz}\tau E^{2}$, where $\tau=\hbar/2\gamma$ is the relaxation time. The coefficient $\beta_{zz}$ is expected to be quantized under appropriate conditions [68, 64, 69]. Here, we evaluate the coefficient $\beta_{zz}$ by fitting the calculated photocurrent $J_{z,0}$ in the weak $E$-field regime to a quadratic curve. The coefficient $\beta_{zz}$ is shown in the inset of the bottom panel. $\beta_{zz}$ changes its sign from negative to positive around $\hbar\Omega=0.3\leavevmode\nobreak\ \mathrm{eV}$, and reaches a maximum near $\hbar\Omega=0.4\leavevmode\nobreak\ \mathrm{eV}$. Under the current conditions, the quantization of $\beta_{zz}$ is not observed, consistent with previous numerical simulations [59, 60]. As the $E$-field increases, higher-order corrections become significant, leading to deviations from the quadratic approximation. Above $E\simeq 0.5\leavevmode\nobreak\ \mathrm{MV/cm}$, $J_{z,0}$ as a function of $\Omega$ exhibits multiple sign changes, where higher-order terms in powers of $E$ play a significant role in determining the sign of $J_{z,0}$. Note that the sign of $J_{z,0}$ can be reversed by changing the handedness of circularly polarized light, reflecting the chirality of CoSi.
@@
Figure 3(a) shows the photocurrent $J_{z,0}$, including the strong $E$-field regime $(|E|\gtrsim 1\leavevmode\nobreak\ \mathrm{MV/cm})$. At relatively low frequencies $\Omega\lesssim 0.5\leavevmode\nobreak\ \mathrm{eV}$, $J_{z,0}$ is positive and increases with the $E$-field, whereas at higher frequencies $\Omega\gtrsim 0.5\leavevmode\nobreak\ \mathrm{eV}$, $J_{z,0}$ becomes negative. In this regime, the nonlinearity becomes more pronounced, and the electronic structure deviates significantly from equilibrium. Figure 3(b) shows the spectrum $N_{{\bf\it k}\omega}$ under the intense $E$-field condition. The spectrum reveal highly asymmetric features, reflecting the hybridization of photon-dressed bands. The spectrum is broadened and widely distributed, further reflecting the strongly driven nonequilibrium nature of the chiral system.
@@
The parity-dependent directivity of HHG can be understood based on dynamical symmetry [70, 71, 72, 73, 74, 75, 76, 77, 78]. From the combination of the screw rotation symmetry of B20-type materials, $s_{2z}=\{C_{2z}|\frac{1}{2}0\frac{1}{2}\}$, and the time periodicity of the photocurrent $J_{\alpha}(t+T)=J_{\alpha}(t)$, we obtain $J_{\parallel}(t+T/2)=e^{-i\pi}J_{\parallel}(t)$ and $J_{\perp}(t+T/2)=J_{\perp}(t)$. These relations result in the following selection rules for HHG: $I_{\parallel}(n)=0$ except for $n\equiv 1\pmod{2}$ and $I_{\perp}(n)=0$ except for $n\equiv 0\pmod{2}$.
To clarify the $E$-field dependence of HHG in more detail, Figure 4(b) shows log-log plots of the HHG intensity as a function of $E$. In the region where the $E$-field is relatively weak ($E\lesssim 1\leavevmode\nobreak\ \mathrm{MV/cm}$), the $n$th order HHG intensity is proportional to $E^{2n}$. This behavior is consistent with the $n$th order response of $J_{\alpha,n}(\propto E^{n})$ in the perturbative regime. As the $E$-field increases, the $E^{2n}$ lines intersect in $E\gtrsim 1\leavevmode\nobreak\ \mathrm{MV/cm}$, indicating the breakdown of perturbative picture. In the strong $E$-field regime, non-perturbative effects become significant. These effects include highly hybridized photon-dressed bands and contributions from high-energy photon-dressed states. As a result, deviations from the $E^{2n}$ lines are observed.
Finally, the frequency $\Omega$ dependence of the HHG intensity is explained. Figure 4(c) shows semilog plots of the HHG intensity as a function of $\Omega$ for the fixed value of $E=2.0\leavevmode\nobreak\ \mathrm{MV/cm}$. For $n=1$, $I_{\parallel}$ increases monotonically with $\Omega$. For $n>2$, the HHG intensity tends to decrease as $\Omega$ increases. Furthermore, as $n$ increases, the reduction in HHG intensity with increasing $\Omega$ becomes more pronounced. This suggests that HHG detection is more feasible at lower frequencies.
IV Discussion
The selection rules discussed above can be extended to systems with an $N$-fold rotational or screw symmetry axis. When light propagates along either of these axes, the in-plane and out-of-plane photocurrents satisfy the following conditions: